A Bibliography of Publications of Ephraim Nissan

Ephraim Nissan

Tel: ?n/a? FAX: ?n/a?

E-mail: E.Nissan@gre.ac.uk (Internet)

13 October 2017 Version 1.12

Abstract

This bibliography records publications of Ephraim Nissan.

Title word cross-reference

12th [PU94]. 1985/June [COG85]. 1987 [OV89]. 1995 [Kel96].

4-7 [COG85].

5th [OV89].

6th [RZ91].

’91 [RZ91].

= [COG85].

A1 [NR97]. Åqvist [NS96b]. acceptations [Nis88a]. account [NW94b, NS96d]. acquisition [Nis87c, Nis91d, Nis95b]. activities [Nis87d]. adaptation [NSGK96]. admissibility [Nis87b]. Advances [Gol90]. advice [AN92]. Adviser [GNS87]. AI [Nis89b, Nis89c, SNG96]. AJIT [Nis08]. Akdeniz [Nis96a]. ALIBI [FEKN^{+}93, KNP91]. alimentation [Nis87d]. allocation [GKN^{+}95, SNG96]. Alternative [KNP91]. Analysis [NS95a, Nis87a, Nis89a, NW94a]. animal [Nis93c, Nis96c, NS96c]. ANN [SNG96]. Appendix [SNG96]. Application [GN88, Nis89b]. application-domains [Nis89b]. Applications [Kel96, TRLA92, Kel96]. approach [Nis92b, SNG96]. April [OV89]. architecture [Nis85b, NSG94, Nis95b]. areas [Nis87d]. argument [NS96d, Nis08]. arguments [NS96e, NS96f]. Artificial [Gol90, Neg91, Nis89b, Nis89c, Nis89a, Nis91a, Nis93b, COG85, Nis91c]. artificiale [Nis93a]. artificielle [COG85]. aspects [Nis91a, Nis95b, NS96a]. assisted [GN89]. assisting [GKN93]. Association [OV89]. attack [GNS95]. attribute [Nis87c]. Attributes [Nis88a, Nis88b]. augmented [Nis95g]. AURANGZEB [Nis08]. automated [SNG96]. automation [NSGK94, NSGK96]. autopoiesis [Nis87c]. Aviv [Pad89].

B [NW94b]. base [Nis88a, Nis88b]. based [Gol90, Nis87d, Nis87e, Nis92b, Nis95b]. bases [Nis87g, Nis88c]. basi [BNP80, Nis82]. BASKETBALL [NSZ92]. behavior [FEKN^{+}93, Nis93b]. Bernardini [Nis89a]. Bible [Nis85a]. bird [SNG96].

calculus [Nis87a, NS96b]. Carolina [RZ87]. CaroMemphis [Nis96c]. case [Nis92c, Nis96c, NS96c, NS96e, NS96f]. case-study [NS96e, NS96f]. census [Nis96a]. Center [Pad89]. Chairs [Nis85c]. Charlotte [RZ87, RZ91]. clarity [Nis87f, Nis87g]. Cognitiva [COG85]. cognitives [COG85]. coinage [Nis85a, Nis85c, Nis85d]. coining [Nis84]. collective [Nis93b]. combining [NSGK94]. come [Nis93a]. common [Nis92a]. commonsense [Nis96c]. community [Nis92c]. compilatore [AFN^{+}81]. computation [NSGK94, Nisxx]. computational [Nis92c]. Computer [NS95a, TRLA92, Nis87d, Nis89b, Nis91b]. concentric [GNS95]. concept [Nis89c]. conception [GNS95]. Conceptual [NS95a]. concettuale [Nis82]. Conference [Pad89, PU94, TRLA92]. configurations [GN88, NSGK96]. Congress [CG87, OV89]. connaissance [COG85]. considerations [NWY96]. considerazioni [BNP80]. Consistency [MMN96]. constructing [Nis88c]. constructions [Nis96c, NS96c]. Content [NS95a]. Convention [Pad89]. coping [Nis92a]. core [GKN93, NSGK96]. Could [Nis85a]. couldn’t [Nis85a]. counterplanning [NSZ92]. courses [Nis89b]. CRWO [OV89]. cultural [Nis92c, Nis96c, NS96c]. CuProS [Nis88c]. customizing [Nis86]. Cyberspace [Nis93c].

Data [Nis87a]. Database [Nis83, GNS87]. dati [BNP80, Nis82]. deception [Nis08]. deductive [GNS87]. deeper [Nis88b]. defensible [KNP91]. definition [Nis86].

Deontic [NS96a]. deontically [KNP91]. Derivational [Nis92a]. Derivative [Nis83]. describing [Nis88c]. Design [Nis83, GKN93]. Development [OV89]. Deviation [Nis92b]. dialectic [Nis91a]. dialettica [Nis93a]. dictionary [Nis90a]. Diglossia [Nis96b]. DINING [GNS87, Nis87h]. discourse [Nis95b]. Discovery [GN89, NSGK94]. dogs [NS95b, NS96c]. domain [GN89, NSG94]. domains [Nis89b]. domande [BNP80]. Dutch [OV89]. dynamics [Nis87f, Nis92c].

EARDHE [OV89]. ecosystem

[NS96e, NS96f]. editing [NW94b]. editions [NW94b]. editor [Nis86]. editorial [NGC08, NGC09]. Education [TRLA92, Nis89b, Nis89c, OV89]. Electrical [Pad89]. Electronics [Pad89]. ELLE [AFN^{+}81]. energy [Kel96]. engineering [CG87, GN89, GKN93, GN96, Nis87e, Nis90a, NSG94, NSGK94, SNG96]. Engineers [Pad89]. entries [Nis88a]. entry [Nisxx]. Environmental [Kel96]. Epistemic [Nis08]. Epistemological [NWY96]. Etruscan [Nis89a, Nis91b, Nis92c]. European [OV89]. evaluation [Nis87f, Nis87g, NSZ92]. events [Nis95g]. evidence [NS96b, NR97]. evidential [NS96b]. evironment [Kel96]. Exception [Nis87b]. Exception-admissibility [Nis87b]. excuses [KNP91]. exegesis [NW94a]. exemplification [Nis93b]. exercise [Nis96c]. Expert [Nis88d, AN92, GN89, GKN93, GN96, Nis84, Nis85b, Nis85a, Nis85c, Nis85d, Nis86, Nis87f, Nis87g, Nis87h, Nis88c, NSZ92]. expression [Nis95a]. eye [SNG96].

Fair [Pad89]. Federal [CG87]. FIDEL [Nis87h]. Finding [KNP91]. food [Nis87h]. food-processing [Nis87h]. Formal [Nis83, NS96d, Nis08]. formali [Nis82].

Formalism [NS95b, NS96c, Nis88c, Nis96c]. formalisms [NR97]. formalizing [NS96e, NS96f]. Formation [Nis88d, Nis87f, Nis87g, Nis90b, Nis91d, Nisxx]. formed [Nis96a]. formula [Nis96a]. formulae [Nis08]. fostering [Nis91c]. fra [Nis93a]. frame [Nis86]. frame-definition [Nis86]. frames [Nis85c, Nis85d, Nis87e]. framework [Nis95b]. frontier [COG85]. frontiére [COG85]. fuel [GN88, GKN93, GNS95, GKN^{+}95, NSGK96, SNG96]. fuel-allocation [SNG96]. FUELCON [GKN93, GNS95, GN96]. full [NSGK94]. Future [Neg91].

GASTRO [Nis87h]. gastronomy [Nis87h]. general [Nis87g]. generalized [wHW96a, wHW96b, wHW96c]. generation [GN88, NSGK96]. Genesis [NW94a]. geometrical [Nis87a]. Germany [CG87]. gist [NS96d]. given [NW94b]. Going [Nis88b]. grammar [Nis95g]. grammars [Nis95f]. Grounds [Pad89]. Guest [NGC08, NGC09].

health [Kel96]. Hebrew [LN91]. held [RZ87]. heuristic [GN88, GN89]. heuristics [NSGK94, NSGK96, SNG96]. Higher [OV89, Nis89b, Nis89c, OV89]. homiletics [NW94a]. hospitality [Nis87d]. host [Nis86]. human [Nis93b, Nis96c, NS96c]. husbandry [Nis93c]. hybrid [GNS95, SNG96]. hybrid-paradigm [GNS95]. Hypergraph [Nis95f, Nis95g]. hypergraph-grammar [Nis95g]. HyperJoseph [NW94a, NW94b, NWY96, wHW96a, wHW96b, wHW96c]. Hypermedia [NW94a]. hypertextual [NWY96].

IAPR [PU94]. identity [Nis95f, Nis08]. IEEE [Pad89]. II [Nis85d, Nis87g, Nis87h, Nis88b, Nis89c, Nis95d, Nis96c, NS96e, NS96f, wHW96b]. III [Nis95e, wHW96c]. imputed [FEKN^{+}93]. in-core [GKN93, NSGK96]. incompleteness [Nis92a]. individual [Nis93b]. Indoglottal [Nis89a]. Info [Nis83]. Info-Spatial [Nis83]. Information [NS95a]. Insights [Nis92c]. instances [Nis96a]. Institute [TRLA92]. integrated [NSG94]. integration [GKN^{+}95]. Intelligence [Gol90, COG85, Nis89b, Nis89c, Nis89a, Nis91a, Nis91c, Nis93b, COG85, Neg91]. intelligent [RZ87, RZ91]. International [CG87, PU94, RZ91, TRLA92, RZ87]. Interpretation [FEKN^{+}93, Nis89a, Nis92c]. interpreting [Nis90b]. intertextuality [NW94b]. inventing [Nis85b]. ISMIS [RZ91]. isolation [Nis95b]. Israel [Pad89, PU94, TRLA92]. issue [NGC08, NGC09]. Issues [Nis90a].

January [TRLA92]. Jerusalem [PU94]. juin [COG85]. June [COG85].

Kappa [NS96b]. Knowledge

[Gol90, Nis87c, Nis87d, NS95a, CG87, Nis87e, Nis87g, Nis88a, Nis88b, Nis88c, Nis89c, Nis89a, Nis92a, Nis92b, NW94a, Nis95b]. knowledge-analysis [Nis89a]. knowledge-base [Nis88a, Nis88b]. Knowledge-based [Gol90, Nis87d, Nis92b, Nis95b]. knowledge-bases [Nis87g, Nis88c]. knowledge-presentation [Nis89c].

Laboratory [Kel96]. lambs [Nis95f].

Language

[Gol90, Nis86, Nis88c, Nis90b, Nis91c, Nisxx]. layer [NSZ92]. layout [NW94b]. learning [NSG94]. Legal [MMN96, Nis95g, NS96b, NR97]. legendary [NW94a]. leopards [Nis85d]. Lexical [Nis92d, Nis85a, Nis88a, Nis91d, Nis92a, Nis95b]. lexicographical [Nis88a, Nis88b]. lexicon [Nis90a]. liberal [Nis88c]. line [GNS95]. linguaggio [AFN^{+}81].

L’intelligenza [Nis93a]. lions [Nis85d]. lodging [Nis87d]. logical [NS96b]. logico [Nis82]. logico-concettuale [Nis82]. loop [GN89]. lost [Nis95f].

machine [Nis84, Nis90a, NSG94]. machine-dictionary [Nis90a]. machine-made [Nis84]. made [Nis84]. management [GKN93, GNS95, Nis87d, NSGK96]. managing [Nis88c]. manipulation [Nis90a]. March [Kel96, Pad89]. Marco [NGC08, NGC09]. Marzolla [Nis89a]. Meanings [Nis95a, Nis95c, Nis95d, Nis95e]. meant [Nis90a]. memorial [NGC08, NGC09]. metamodel [Nis89a, Nis92c]. metarepresentation [Nis87c, Nis87e]. metatheory [Nis89a]. method [GN88]. Methodologies [RZ87, RZ91]. Mint [Nis92d, Nis85a]. misantonymously [Nis96a]. model [Nis87f]. modelling [NS96e, NS96f]. models [Nis91b, Nis92b, Nis93b]. Monographic [Nisxx]. morpho [Nis87f, Nis87g]. morpho-semantic [Nis87f, Nis87g]. Moses [Nis85a]. multifaceted [Nis93b]. multilingual [Nis90a]. multiple [NSZ92]. multiple-layer [NSZ92]. mutualism [NS96e, NS96f].

N.C [RZ91]. names [Nis96a]. narrative [NW94a, Nis08]. Natania [TRLA92]. National [Kel96, Nis91c]. native [Nis91c]. Natural [Gol90]. nel [Nis82]. Neologisms [Nis95b, Nis85b]. Nested [Nis87e, Nis88c]. Nested-relation [Nis87e, Nis88c]. network [NSG94]. networks [Kel96]. neural [Kel96, NSG94, NSGK96, SNG96]. neural/symbolic [SNG96]. neurosciences [COG85]. NLP [LN91]. North [RZ87]. Northwest [Kel96]. note [NS96b]. novel [SNG96]. nuclear [GKN93, GNS95, GKN^{+}95, GN96, NSG94, NSGK94, NSGK96, SNG96]. nuclear-fuel [GNS95].

Oct [CG87]. October [PU94, RZ87, RZ91]. on-line [GNS95]. ongoing [GNS95]. ONOMATURGE [Nis84, Nis85b, Nis85a, Nis85c, Nis85d, Nis87f, Nis87g, Nis88a, Nis88b, Nis88d, Nis91c]. optimal [GKN93]. optimized [SNG96]. organization [Nis88a, NWY96]. other [Nis91a]. outlined [GNS95]. Overall [Nis88a].

Pacific [Kel96]. Packaging [Nis96b]. page [NW94b]. papers [Pad89]. paradigm [GNS95, Nis91c]. Paratextual [Nis96b]. Paris [COG85]. Part [GNS87, Nis85c, Nis85d, Nis87f, Nis87g, Nis87h, Nis88a, Nis88b, Nis89b, Nis89c, NW94a, NW94b, Nis95c, Nis95d, Nis95e, Nis96c, NS96c, NS96e, NS96f, wHW96a, wHW96b, wHW96c]. partire [BNP80]. partonomies [NS95b, NS96c]. paschal [Nis95f]. Pattern [PU94]. people [Nis84]. perform [Nis85a]. perspective [Nis87f]. Physical [TRLA92]. plans [KNP91]. policies [Nis91c]. power [Nis96c]. practice [GKN93]. PragMathematics [Nis96c, NS96c, NS96a]. pragmatics [Nis93b, Nis96a]. predicate [Nis87a]. prediction [GKN^{+}95]. Predictive [AN92]. prescriptions [Nis95g]. Present [Nis89c]. presentation [Nis89c]. problem [GNS95, GN96]. Proceedings [CG87, OV89, RZ87, Kel96, PU94, TRLA92, RZ91]. processing [Nis87h]. Progetto [BNP80, Nis82]. program [Pad89]. Project [GNS87, Nis84, Nis87h, NW94a, NW94b, GN96, NSGK94, NS96a]. projects [GNS95]. Proprietà [Nis82]. proto [Nis87b]. proto-representations [Nis87b]. prototype [Nis96a]. prototype-theoretic [Nis96a]. prototypes [Nis95a, Nis95c, Nis95d, Nis95e].

RAFFAELLO [Nis86, Nis87e]. reactors [GKN^{+}95]. Recognition [PU94]. recreational [Nis87d]. refinement [GN89]. regulation [Nis92b]. related [Nis85c, Nis85d, Nis87d]. relation [Nis87e, Nis88c]. reload [GN88]. repertoire [Nis95c, Nis95d, Nis95e]. Representation [Nis87e, Nis85c, Nis85d, Nis87a, Nis87g, Nis90a, NW94a, NW94b, Nis95g, Nis08]. representations [Nis87b, Nis96c]. Republic [CG87]. rescue [Nis95f]. Research [OV89, GKN93, Nis92c]. Resource [NSZ92]. respective [NW94a]. revision [GKN^{+}95, SNG96]. Richland [Kel96]. role [Nis96c, NS96c]. rule [GKN^{+}95]. rule-revision [GKN^{+}95]. rules [GN89]. rulesets [NSZ92].

Schoenberg [Nis85a]. scholarly [NS96d]. science [Nis89b, Nis91a]. sciences [COG85]. scienza [Nis93a]. search [GN88]. second [RZ87]. semantic [Nis87f, Nis87g]. semantics [Nis87e]. Semitic [Nis90a]. sense [Nis92a]. SEPPHORIS [Nis95g]. Sept [CG87]. sequel [GNS95]. Shape [Neg91]. Sixteenth [Pad89]. SKILL [FEKN^{+}93]. SKINDIVING [AN92]. Somalvico [NGC08, NGC09]. sources [Nis89c]. Spatial [Nis83]. sperimentale [AFN^{+}81]. Sport [TRLA92]. Statements [MMN96]. statics [Nis87g]. stipulations [Nis95g]. strength [NS96b]. Structure [Nis90a, Nis86, Nis87e]. structure-editor [Nis86]. structures [Nis88c, Nis08]. study [NS96e, NS96f]. style [Nis95b]. subproject [Nis08]. subsymbolic [GKN^{+}95, NSGK94]. sulle [BNP80]. supporting [Nis91c, NW94a]. Symbolic [GKN^{+}95, Nis90a, NSG94, NSGK94, NSGK96, SNG96]. Symposium [RZ91, RZ87]. synonyms [Nis85c, Nis85d]. syntax [NW94b]. System [Nis88d, AN92, GKN93, GN96, Nis84, Nis85b, Nis85a, Nis85c, Nis85d, Nis87f, Nis87g, Nis87h, NSZ92]. systems [Nis86, Nis87d, RZ87, RZ91, Gol90].

takes [NW94b]. TAMBALACOQUE [NS96d, NS96e, NS96f]. task [Nis87f]. tasks [GKN^{+}95, Nis87d]. technologies [OV89]. technology [Nis91a]. tecnologia [Nis93a]. Tel [Pad89]. Tel-Aviv [Pad89]. Temporal [MMN96]. terminal [Nis87h]. Terminology [CG87, Nis87g, CG87]. terms [Nis85c, Nis85d, Nis95b]. text [NW94a, Nis95b]. theoretic [Nis96a]. Theory [Nis90b, NS96b, wHW96a, wHW96b, wHW96c]. tongue [Nis85a]. Tool [Nis83, GN89, Nis88c, Nis91c]. tools [SNG96]. Topics [Nis84]. topology [Nis88b]. transition [NS95b, NS96c]. translation [NW94b]. transmitted [Nis96c]. trends [Nis89c]. trials [Nis92c]. triangles [Nis96c]. Trier [CG87]. Twelve [Nis85c]. typicality [Nis87b].

Understanding [Neg91]. University [CG87]. unknown [Nis90b]. Upgrading [NSGK96]. USA [Kel96, RZ91]. use [Nis84]. using [Nis87a].

VegeDog [NS95b, NS96c]. vegetarian [NS95b, NS96c]. view [SNG96]. viewpoint [Nis87g]. vs [Nis96c, NS96c].

Washington [Kel96]. Wine [GNS87]. Wingate [TRLA92]. WINING [GNS87, Nis87h]. within [Nis08]. Word [Nis85a, Nis88d, Nis91d, Nisxx, Nis84, Nis85c, Nis85d, Nis87f, Nis87g]. Word-coinage [Nis85a, Nis85c, Nis85d]. word-coining [Nis84]. Word-Formation [Nis88d, Nis91d, Nis87f, Nis87g]. words [Nis84]. Workshop [Kel96]. Would [Nis84].

References

Abbati:1981:CSP

[AFN^{+}81] Paolo Abbati, Alfonso Fuggetta, Ephraim Nissan, Alfredo Sartorato, and Vincenzo Viscuso. Un compilatore sperimentale per il linguaggio ELLE. Milan, Italy, 1981. In Italian.

Aviel:1992:PAS

[AN92] Amit Aviel and Ephraim Nissan. Predictive advice in the SKINDIVING expert system. In Tenenbaum et al. [TRLA92], pages 39–44. ISBN ???? LCCN GV436 .I576 1992.

Bozzoli:1979:PBD

[BNP80] Lucio Bozzoli, Ephraim Nissan, and Barbara Pernici. Progetto di basi di dati a partire da considerazioni sulle domande. Tesi di Laurea in Ingegneria Elettronica, Istituto di Elettrotecnica ed Elettronica, Politecnico di Milano (Milan Institute of Technology), Milan, Italy, 1979–1980. (Italian: Database Design Based on Considerations About Queries), p. 235.

Czap:1987:TKE

[CG87] Hans Czap and Christian Galinski, editors. Terminology and knowledge engineering: Proceedings, International Congress on Terminology and Knowledge Engineering, 29 Sept.–1 Oct. 1987, University of Trier, Federal Republic of Germany. Indeks Verlag, Frankfurt/M, Germany, 1987. ISBN 3-88672-202-3 (vol.

1), 3-88672-203-1 (vol. 2). LCCN P305 .I56 1987. Two volumes.

Anonymous:1985:CPJ

[COG85] Cognitiva 85, Paris 4-7 juin 1985/June 4–7, 1985: A la frontiére de l’intélligence artificielle des sciences de la connaissance des neurosciences = At the frontier of artificial intelligence, cognitives sciences, neurosciences. CESTA, Paris, France, 1985. ISBN ???? LCCN ????

Fakher-Eldeen:1993:IIB

[FEKN^{+}93] Fakher Fakher-Eldeen, Tsvi Kuflik, Ephraim Nissan, Gilad Puni, Ronnie Salfati, Yuval Shaul, and Auni Spanioli. Interpretation of imputed behavior in ALIBI (1 to 3) and SKILL. Informatica e diritto, II(1/2):213–242, 1993. ISSN 0390-0975. Year 19, 2nd Series.

Galperin:1993:FES

[GKN93] Alex Galperin, Yehoshua Shuky Kimhy, and Ephraim Nissan. FU-ELCON: an expert system for assisting the practice and research of in-core fuel management and optimal design in nuclear engineering. Computers and Artificial Intelligence = Vychislitel’nye mashiny i iskusstvennyi intellekt, 12(4):369–415, 1993. CODEN CARIDY. ISSN 0232-0274.

Galperin:1995:SSI

[GKN^{+}95] Alex Galperin, Shuky Yehoshua Kimhi, Ephraim Nissan, Hava Siegelmann, and Jun Zhao. Symbolic and subsymbolic integration in prediction and rulerevision tasks for fuel allocation in nuclear reactors. In Proceedings of the Third European Congress on Intelligent Techniques and Soft Computing (EUFIT’95), Vol. 3, pages 1546– 1550. ????, Aachen, Germany, August 28–31, 1995.

Galperin:1988:AHS

[GN88] Alex Galperin and Ephraim Nissan. Application of a heuristic search method for generation of fuel reload configurations. Nuclear Science and Engineering, 99 (4):343–352, 1988. CODEN NSE-NAO. ISSN 0029-5639.

Galperin:1989:DAE

[GN89] Alex Galperin and Ephraim Nissan. Discovery as assisted by an expert tool: a refinement loop for heuristic rules in an engineering domain. In Padan [Pad89], page various. ISBN ???? LCCN ???? Paper 1.4.3.

Galperin:1996:FPN

[GN96] Alex Galperin and Ephraim Nissan. The FUELCON project in nuclear engineering: from the problem to the expert system. Technical Report Report (c/o E. Nissan), CMS, Univ. of Greenwich, Woolwich, London, UK, 1996.

Garber:1987:WDP

[GNS87] Sammy Garber, Ephraim Nissan, and Amnon Shur. The WIN-ING AND DINING project. part I: ‘the wine adviser’, a deductive database. International Journal of Hospitality Management, 6(4):

203–205, 1987. CODEN IJH-MDN. ISSN 0278-4319. Special Issue on Information Technology.

Galperin:1995:HPC

[GNS95] Alex Galperin, Ephraim Nissan, and Hava Siegelmann. The hybrid-paradigm conception of ongoing or outlined sequel projects to FUELCON: a concentric attack on the nuclearfuel on-line management problem. In Keller [Kel96], page ???? ISBN 981-02-2758-2. LCCN QA76.87.W675 1995.

Golumbic:1990:AAI

[Gol90] Martin Charles Golumbic, editor. Advances in Artificial Intelligence: Natural Language and Knowledge-based Systems. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1990. ISBN 0-387-97355-9 (New York), 3-540-97355-9 (Berlin). LCCN QA76.9.N38 A34 1990. Based on papers from three conferences held in Israel 1987-1989.

Keller:1996:ANN

[Kel96] Paul E. Keller, editor. Applications of neural networks in evironment, energy, and health: proceedings of the 1995 Workshop on Environmental and Energy Applications of Neural Networks, Pacific Northwest National Laboratory, Richland, Washington, USA, 30–31 March 1995. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1996. ISBN 981-02-2758-2. LCCN QA76.87.W675 1995.

Kuflik:1991:FEA

[KNP91] Tsvi Kuflik, Ephraim Nissan, and Gilad Puni. Finding excuses with ALIBI: Alternative plans that are deontically more defensible. Computers and Artificial Intelligence = Vychislitel’nye mashiny i iskusstvennyi intellekt, 10(4):297–325, 1991. CODEN CARIDY. ISSN 0232-0274.

Lancashire:1991:H

[LN91] Ian Lancashire and Ephraim Nissan. Hebrew [NLP]. In Ian Lancashire, editor, The Humanities Computing Yearbook, 1989–90, pages 292–300. Clarendon Press, Oxford, UK, 1991.

Ma:1996:TCL

[MMN96] Brian Ma, Jixin Ma, and Ephraim Nissan. Temporal consistency of legal statements. Technical Report Report (c/o E. Nissan), CMS, Univ. of Greenwich, Woolwich, London, UK, 1996.

Negrotti:1991:UAF

[Neg91] Massimo Negrotti, editor. Understanding the Artificial. On the Future Shape of Artificial Intelligence. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1991. ISBN 3-540-19612-9 (Berlin), 0-387-19612-9 (New York). LCCN Q335 .N435 1991.

Nissan:2008:GEM

[NGC08] Ephraim Nissan, Giuseppina Gini, and Marco Colombetti. Guest editorial: Marco Somalvico memorial issue. Annals of Mathematicsl and Artificial Intelligence, 54(4):257–264 (2009), 2008. CODEN AMAIEC. ISSN 1012-2443.

Nissan:2009:GEM

[NGC09] Ephraim Nissan, Giuseppina Gini, and Marco Colombetti. Guest editorial: Marco Somalvico memorial issue. Comput. Intell., 25(2):109–113, 2009. ISSN 0824-7935.

Nissan:1982:PFN

[Nis82]  Ephraim Nissan.  Proprietà

formali nel progetto logicoconcettuale di basi di dati. Tesi di Laurea in Ingegneria Elettronica, Dipartimento di Elettronica, Politecnico di Milano (Milan Institute of Technology), Milan, Italy, 1982. (Italian: Formal Properties in the Logical and Conceptual Design of Databases), 2 vols., p. 400+200. Awarded the Burroughs Italiana Prize.

Nissan:1983:ISD

[Nis83] Ephraim Nissan. The info-spatial derivative: A new formal tool for database design. In Proceedings of the AICA’83 Conference, volume 2, pages 177–182. Associazione Italiana per il Calcolo Automatico (Milan), Naples, Italy, September 28–30, 1983.

Nissan:1984:WPU

[Nis84] Ephraim Nissan. Would people use machine-made words? (and why?) A word-coining expert system: Topics in the ONOMA-TURGE project. Technical report, Report available from the author, 1984.

Nissan:1985:CES

[Nis85a] Ephraim Nissan. Could an expert system perform what schoenberg couldn’t for moses? word-coinage in the bible’s tongue: ONO-MATURGE, a lexical mint. In COGNITIVA-85 [COG85], pages 95–100. ISBN ???? LCCN ????

Nissan:1985:AOE

[Nis85b] Ephraim Nissan. On the architecture of ONOMATURGE, an expert system inventing neologisms. In Proceedings of the 12th Conference of the Association for Literary and Linguistic Computing, volume 2, pages 671–680. Champion-Slatkine, Geneva and Paris, Nice, France, 1985.

Nissan:1985:RSRa

[Nis85c] Ephraim Nissan. The representation of synonyms and of related terms, in the frames of an expert system for word-coinage. part I: The twelve chairs and ONOMA-TURGE. In Proceedings of the Fifth International Workshop on Expert Systems and Their Applications, volume 2, pages 685– 703. ????, Avignon, France, June 1985.

Nissan:1985:RSRb

[Nis85d] Ephraim Nissan. The representation of synonyms and of related terms, in the frames of an expert system for word-coinage. part II: On lions, leopards and ONOMA-TURGE. In Proceedings of the Fifth International Workshop on Expert Systems and Their Applications, volume 2, pages 705– 741. ????, Avignon, France, June 1985.

Nissan:1986:FDL

[Nis86] Ephraim Nissan. The framedefinition language for customizing the RAFFAELLO structureeditor in host expert systems. In Zbigniew Ras and Maria Zemankova, editors, Proceedings of the First International Symposium on Methodologies for Intelligent Systems (ISMIS’86), pages 8–18. ACM SIGART Press, New York, Knoxville, TN, USA, 1986.

Nissan:1987:DAU

[Nis87a] Ephraim Nissan. Data analysis using a geometrical representation of predicate calculus. Information Sciences, 41(3):187– 258, 1987. CODEN ISIJBC. ISSN 0020-0255 (print), 1872-6291 (electronic).

Nissan:1987:EAT

[Nis87b] Ephraim Nissan. Exceptionadmissibility and typicality in proto-representations. In Czap and Galinski [CG87], pages 253– 267. ISBN 3-88672-202-3 (vol. 1), 3-88672-203-1 (vol. 2). LCCN P305 .I56 1987. Two volumes.

Nissan:1987:KAM

[Nis87c] Ephraim Nissan. Knowledge acquisition and metarepresentation: attribute autopoiesis. In Ras and Zemankova [RZ87], pages 240–247. ISBN 0-444-01295-8. LCCN Q334 .I576 1987.

Nissan:1987:KBC

[Nis87d] Ephraim Nissan. Knowledgebased computer systems for tasks in hospitality management or related areas: recreational activities, alimentation, and lodging. International Journal of Hospitality Management, 6(4):199–202, 1987. CODEN IJHMDN. ISSN 0278-4319. Special Issue on Information Technology.

Nissan:1987:NRB

[Nis87e] Ephraim Nissan. Nested-relation based frames in RAFFAELLO. representation and metarepresentation structure and semantics for knowledge engineering. In International Workshop on Theory and Applications of Nested Relations and Complex Objects, pages 95–99. INRIA, France, Darmstadt, Germany, 1987.

Nissan:1987:OESa

[Nis87f]  Ephraim Nissan.  ONOMA-

TURGE: An expert system for word-formation and morphosemantic clarity evaluation. part I: The task in perspective, and a model of the dynamics of the system. In Czap and Galinski [CG87], pages 167–176. ISBN 3-88672-202-3 (vol. 1), 3-88672-203-1 (vol. 2). LCCN P305 .I56 1987. Two volumes.

Nissan:1987:OESb

[Nis87g]  Ephraim Nissan.  ONOMA-

TURGE: An expert system for word-formation and morphosemantic clarity evaluation. part II: The statics of the system. the representation from the general viewpoint of knowledge-bases for terminology. In Czap and Galinski [CG87], pages 177–189. ISBN 3-88672-202-3 (vol. 1), 3-88672-203-1 (vol. 2). LCCN P305 .I56 1987. Two volumes.

Nissan:1987:WDP

[Nis87h] Ephraim Nissan. The WINING AND DINING project. part II: FIDEL GASTRO, an expert system for gastronomy and terminal food-processing. International Journal of Hospitality Management, 6(4):207–215, 1987. CO-DEN IJHMDN. ISSN 0278-4319. Special Issue on Information Technology.

Nissan:1988:ALKa

[Nis88a] Ephraim Nissan. Attributes in the lexicographical knowledgebase of ONOMATURGE. part I: Overall organization of lexical entries and acceptations. Technical Report Report (available from the author), CMS, Univ. of Greenwich, Woolwich, London, UK, 1988.

Nissan:1988:ALKb

[Nis88b] Ephraim Nissan. Attributes in the lexicographical knowledgebase of ONOMATURGE. part II: Going deeper inside the topology. Technical Report Report (available from the author), CMS, Univ. of Greenwich, Woolwich, London, UK, 1988.

Nissan:1988:CLL

[Nis88c] Ephraim Nissan. The CuProS language: a liberal formalism for describing nested-relation structures to an expert tool for constructing and managing knowledge-bases. Technical report, Report available from the author. (Also, appendix of PhD dissertation on ONOMA-TURGE), 1988.

Nissan:1988:OES

[Nis88d]  Ephraim Nissan.  ONOMA-

TURGE: An Expert System in Word-Formation. PhD thesis, Department of Mathematics and Computer Science, Ben-Gurion University of the Negev, Beer-Sheva, Israel, 1988. 600 pp. (3 vols.). In English, with Hebrew abstract. Project awarded the 1988 IPA Award in Computer Science, of the Information Processing Association of Israel.

Nissan:1989:AIM

[Nis89a] Ephraim Nissan. Artificial intelligence for a metatheory of interpretation. (A knowledge-analysis of bernardini marzolla’s indoglottal interpretation of Etruscan: For a metamodel of interpretation). Technical Report Report, 133 p. (available from the author), CMS, Univ. of Greenwich, Woolwich, London, UK, 1989.

Nissan:1989:AIHb

[Nis89b] Ephraim Nissan. Artificial intelligence in higher education. part I: Education for AI. from courses in computer science to courses in application-domains. In Oosthoek and Vroeijenstijn [OV89], pages 249–261. ISBN 0-08-037261-9. LCCN LB1028.3 .E961 1987. US$66.00.

Nissan:1989:AIHc

[Nis89c] Ephraim Nissan. Artificial intelligence in higher education. part II: AI for education. present trends, as sources for a concept of knowledge-presentation. In Oosthoek and Vroeijenstijn [OV89], pages 67–98. ISBN 0-08-037261-9. LCCN LB1028.3 .E961 1987. US$66.00.

Nissan:1990:SRM

[Nis90a] Ephraim Nissan. Structure and representation for a multilingual semitic machine-dictionary: Issues in engineering a lexicon meant for symbolic manipulation. Technical Report Report (available from the author), CMS, Univ. of Greenwich, Woolwich, London, UK, 1990.

Nissan:1990:TFI

[Nis90b] Ephraim Nissan. Theory formation for interpreting an unknown language. In Golumbic [Gol90], pages 95–113. ISBN 0-387-97355-9 (New York), 3-540-97355-9 (Berlin). LCCN QA76.9.N38 A34 1990. Based on papers from three conferences held in Israel 1987-1989.

Nissan:1991:AID

[Nis91a] Ephraim Nissan. Artificial intelligence as a dialectic of science and technology, and other aspects. In Negrotti [Neg91], pages 77–90. ISBN 3-540-19612-9 (Berlin), 0-387-19612-9 (New York). LCCN Q335 .N435 1991. Also published in the Italian version of the volume.

Nissan:1991:EMF

[Nis91b] Ephraim Nissan. Etruscan [computer models for]. In Ian Lancashire, editor, The Humanities Computing Yearbook, 1989–90, page 246. Clarendon Press, Oxford, UK, 1991.

Nissan:1991:OAI

[Nis91c]  Ephraim Nissan.  ONOMA-

TURGE: An artificial intelligence tool and paradigm for supporting national and native language fostering policies. AI and Society, 5 (3):202–217, 1991. CODEN AIS-CEM. ISSN 0951-5666.

Nissan:1991:WFL

[Nis91d] Ephraim Nissan. Word-formation and lexical acquisition. In Proceedings of the Workshop on Natural Language Learning (W.14), at the 12th International Joint Conference on Artificial Intelligence, pages 58–62. ????, Sydney, Australia, August 24–30, 1991.

Nissan:1992:DKC

[Nis92a] Ephraim Nissan. Derivational knowledge and the common sense of coping with the incompleteness of lexical knowledge. In João Lopes Alves, editor, Information Technology and Society: Theory, Uses, Impacts [a selection from the Proceedings of the International Symposium on Communication, Meaning and Knowledge vs. Information Technology, Lisbon, September 13– 15, 1989 (4 vols.)], pages 462– 477. Associação Portuguesa para o Desenvolvimento das Comunicações (APDC), and Sociedade Portuguesa de Filosofia (SPF), Lisbon, 1992.

Nissan:1992:DMR

[Nis92b] Ephraim Nissan. Deviation models of regulation: A knowledgebased approach. Informatica e diritto, I(1/2):181–212, 1992. CODEN ???? ISSN 0390-0975. Year 18, 2nd Series.

Nissan:1992:CMI

[Nis92c] Ephraim Nissan. For a computational metamodel of interpretation trials: The case of Etruscan. insights into the cultural dynamics of a research community. Computers and Artificial Intelligence = Vychislitel’nye mashiny i iskusstvennyi intellekt, 11(2):559–583, 1992. CODEN CARIDY. ISSN 0232-0274.

Nissan:1992:LM

[Nis92d] Ephraim Nissan. The lexical mint. Hebrew Linguistics: A Journal for Hebrew Descriptive, Computational, and Applied Linguistics, 36:39–49, 1992. CODEN HELIEK. ISSN 0334-3472. In Hebrew.

Nissan:1990:LAC

[Nis93a] Ephraim Nissan. L’intelligenza artificiale come dialettica fra scienza e tecnologia. In Massimo Negrotti, editor, Capire l’artificiale (1990), Capire l’artificiale: dall’analogia all’integrazione uomomacchina (1993), pages 119–140.

Bollati-Boringhieri, Torino, Italy, 1990–1993. In Italian.

Nissan:1993:AIM

[Nis93b] Ephraim Nissan. Artificial intelligence models for the pragmatics of individual or collective human behavior: A multifaceted exemplification. In Second International Conference on ‘The Culture of the Artificial’. ????, Ascona, Switzerland, 1993.

Nissan:1993:CAH

[Nis93c] Ephraim Nissan. Cyberspace for animal husbandry. Computers and Electronics in Agriculture, 8(3):251–259, 1993. CO-DEN CEAGE6. ISSN 0168-1699.

Nissan:1995:MEP

[Nis95a] Ephraim Nissan. Meanings, expression, and prototypes. Pragmatics and Cognition, 3(2):317– 364, 1995. CODEN PCOGEZ. ISSN 0929-0907.

Nissan:1995:NFL

[Nis95b] Ephraim Nissan. Neologisms in the framework of lexical acquisition: A knowledge-based architecture, for terms in isolation, and for discourse and style aspects as in text. Technical Report Report (available from the author), CMS, Univ. of Greenwich, Woolwich, London, UK, 1995.

Nissan:1995:MPRa

[Nis95c] Ephraim Nissan. On meanings and prototypes: The repertoire. part I. Technical Report Report, ca. 70 p. (available from the author), CMS, Univ. of Greenwich, Woolwich, London, UK, 1995.

Nissan:1995:MPRb

[Nis95d] Ephraim Nissan. On meanings and prototypes: The repertoire. part II. Technical Report Report, ca. 50 p. (available from the author), CMS, Univ. of Greenwich, Woolwich, London, UK, 1995.

Nissan:1995:MPRc

[Nis95e] Ephraim Nissan. On meanings and prototypes: The repertoire. part III. Technical Report Report, ca. 50 p. (available from the author), CMS, Univ. of Greenwich, Woolwich, London, UK, 1995.

Nissan:1995:LIP

[Nis95f] Ephraim Nissan. On the lost identity of paschal lambs: Hypergraph grammars at the rescue. In Ely Merzbach and Moshe Koppel, editors, ????, volume 3, pages 132–150. Aluma (Jerusalem) for Bar-Ilan University/ Math-CS, Ramat-Gan, Israel, 1995. In Hebrew.

Nissan:1995:SAH

[Nis95g]  Ephraim Nissan.  SEPPHO-

RIS: An augmented hypergraphgrammar representation for events, stipulations, and legal prescriptions. Law, Computers, and Artificial Intelligence, 4(1): 33–77, 1995. CODEN LCAIET.

Nissan:1996:APM

[Nis96a] Ephraim Nissan. Akdeniz: The pragmatics of misantonymously formed names. the formula and a prototype-theoretic census of instances. Technical Report Report (available from the author), CMS, Univ. of Greenwich, Woolwich, London, UK, 1996.

Nissan:1996:PPD

[Nis96b] Ephraim Nissan. Paratextual packaging and diglossia. Technical Report Report (available from the author), CMS, Univ. of Greenwich, Woolwich, London, UK, 1996.

Nissan:1996:PIC

[Nis96c] Ephraim Nissan. Pragmathematics, II. caromemphis: An exercise in cultural commonsense on power triangles and transmitted representations. [part II of ‘pragmathematics: formalism and cultural constructions. the case of human vs. animal roles’]. Technical Report Report (c/o E. Nissan), CMS, Univ. of Greenwich, Woolwich, London, UK, 1996.

Nissan:19xx:WFL

[Nisxx] Ephraim Nissan. Word formation in language and computation. [monographic entry]. Encyclopedia of Library and Information Science, 19xx.

Nissan:2008:EFA

[Nis08] Ephraim Nissan. Epistemic formulae, argument structures, and a narrative on identity and deception: a formal representation from the AJIT subproject within AURANGZEB. Annals of Mathematicsl and Artificial Intelligence, 54(4):293–362 (2009), 2008. CODEN AMAIEC. ISSN 1012-2443.

Nissan:1997:TAF

[NR97] E. Nissan and D. Rousseau. Towards A1 formalisms for legal evidence. Lecture Notes in Computer Science, 1325:328–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nissan:1995:IKC

[NS95a] Ephraim Nissan and Klaus M. Schmidt, editors. From Information to Knowledge: Conceptual and Content Analysis by Computer. Intellect Books, Oxford, UK, 1995. ISBN 1-871516-50-1. x + 319 pp. LCCN P98 .F73 1995.

Nissan:1995:VFV

[NS95b] Ephraim Nissan and Solomon Eyal Shimony. Vegedog: Formalism, vegetarian dogs, and partonomies in transition. Technical Report Report, ca. 30 p. (c/o E. Nissan), CMS, Univ. of Greenwich, Woolwich, London, UK, 1995.

Nissan:1996:DAP

[NS96a] Ephraim Nissan and Solomon Eyal Shimony. Deontic aspects in the PragMathematics Project. In Fifth Ibero–American Conference on Computing and Law. ????, ????, 1996.

Nissan:1996:KCE

[NS96b] Ephraim Nissan and Solomon Eyal Shimony. Kappa calculus and evidential strength: A note on åqvist’s logical theory of legal evidence. Technical Report Report (c/o E. Nissan), CMS, Univ. of Greenwich, Woolwich, London, UK, 1996.

Nissan:1996:PVF

[NS96c] Ephraim Nissan and Solomon Eyal Shimony. Pragmathematics, I. vegedog: Formalism, vegetarian dogs, and partonomies in transition. [part I of ‘pragmathematics: formalism and cultural constructions. the case of human vs. animal roles’]. Technical Report Report, ca. 75 p. (c/o E. Nissan), CMS, Univ. of Greenwich, Woolwich, London, UK, 1996.

Nissan:1996:TFA

[NS96d] Ephraim Nissan and Solomon Eyal Shimony. TAMBALACOQUE: For a formal account of the gist of a scholarly argument. Knowledge Organization : KO, 23(3), 1996. CODEN KNOREM. ISSN 0943-7444.

Nissan:1996:TTFa

[NS96e] Ephraim Nissan and Solomon Eyal Shimony. TAMBALACOQUE: Toward formalizing arguments in ecosystem modelling. part II: A case-study in mutualism. Technical Report Report (c/o E. Nissan), CMS, Univ. of Greenwich, Woolwich, London, UK, 1996.

Nissan:1996:TTFb

[NS96f] Ephraim Nissan and Solomon Eyal Shimony. TAMBALACOQUE: Toward formalizing arguments in ecosystem modelling. part II: A case-study in mutualism. Technical Report Report (c/o E. Nissan), CMS, Univ. of Greenwich, Woolwich, London, UK, 1996.

Nissan:1994:ISN

[NSG94] Ephraim Nissan, Hava Siegelmann, and Alex Galperin. An integrated symbolic and neural network architecture for machine learning in the domain of nuclear engineering. In Peleg and Ullman [PU94], pages 494–496. ISBN 0-8186-6265-4 (vol. 1), 0-8186-6270-0 (vol. 2), 0-8186-6275-1 (vol. 3). LCCN TK 7882 P3 I57 1994. Vol. 2: Conference B: Pattern Recognition and Neural Networks.

Nissan:1994:TFA

[NSGK94] Ephraim Nissan, Hava Siegelmann, Alex Galperin, and Shuky Yehoshua Kimhi. Towards full automation of the discovery of heuristics in a nuclear engineering project, by combining symbolic and subsymbolic computation. In Ras and Zemankova [RZ91], pages 427–436. ISBN 3-540-54563-8 (Berlin), 0-387-54563-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 I582 1991.

Nissan:1996:UAN

[NSGK96] Ephraim Nissan, Hava Siegelmann, Alex Galperin, and Shuky Yehoshua Kimhi. Upgrading automation for nuclear fuel in-core management: from the symbolic generation of configurations, to the neural adaptation of heuristics. Engineering with Computers, 1996. CODEN EN-GCE7. ISSN 0177-0667.

Nissan:1992:RECb

[NSZ92] Ephraim Nissan, David Simhon, and Nissim Zigdon. Resource evaluation and counterplanning with multiple-layer rulesets, in the BASKETBALL expert system. In Tenenbaum et al. [TRLA92], pages 60–80. ISBN ???? LCCN GV436 .I576 1992.

Nissan:1994:HPPa

[NW94a] Ephraim Nissan and Hillel Weiss.

The hyperjoseph project. part A: Hypermedia and knowledge representation for supporting an analysis of the text and narrative in genesis 39 and of the respective exegesis and legendary homiletics. In Ferdinand Poswick, editor, Proceedings of the Fourth International Conference on Bible and Computers (AIBI’94), Amsterdam, pages 154–162. Association Internationale Bible et Informatique, Maredsous, Anhee (Denee), Belgium, August 15–18, 1994.

Nissan:1994:HPPb

[NW94b] Ephraim Nissan and Hillel Weiss.

The hyperjoseph project. part B: A representation syntax for intertextuality, that takes into account translation, editing, and the page layout of given editions. In Ferdinand Poswick, editor, Proceedings of the Fourth International Conference on Bible and Computers (AIBI’94), Amsterdam, pages 163–173. Association Internationale Bible et Informatique, Maredsous, Anhee (Denee), Belgium, August 15–18, 1994.

Nissan:1996:HHO

[NWY96] Ephraim Nissan, Hillel Weiss, and Avraham Yossef. Hyperjoseph: The hypertextual organization. epistemological considerations. Knowledge Organization : KO, 23(1):16–24, 1996. CODEN KNOREM. ISSN 0943-7444.

Oosthoek:1989:HEN

[OV89] Hans Oosthoek and Ton Vroeijenstijn, editors. Higher education and new technologies: Proceedings of the 5th Congress of the European Association for Research and Development (EARDHE) and the Dutch Association for Research and Development in Higher Education (CRWO), April 22–25, 1987. Pergamon, New York, 1989. ISBN 0-08-037261-9. LCCN LB1028.3 .E961 1987. US$66.00.

Padan:1989:ISC

[Pad89] Uzi Padan, editor. IEEE, the Sixteenth Conference of Electrical and Electronics Engineers in Israel: program and papers, March 7–9, 1989, Convention Center, Tel-Aviv Fair Grounds. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. ISBN ???? LCCN ???? IEEE Catalog Number 89-CH2723-5.

Peleg:1994:PII

[PU94] Shmuel Peleg and Shimon Ullman, editors. Proceedings of the 12th IAPR International Conference on Pattern Recognition: October 9–13, 1994, Jerusalem, Israel. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6265-4 (vol. 1), 0-8186-6270-0 (vol. 2), 0-8186-6275-1 (vol. 3). LCCN TK 7882 P3 I57 1994. Three volumes.

Ras:1987:MIS

[RZ87] Zbigniew W. Ras and Maria Zemankova, editors. Methodologies for intelligent systems: Proceedings of the second international symposium, held October 14–17, 1987, in Charlotte, North Carolina. North-Holland Publishing Co., Amsterdam, The Netherlands, 1987. ISBN 0-444-01295-8. LCCN Q334 .I576 1987.

Ras:1991:MIS

[RZ91] Zbigniew W. Ras and Maria Zemankova, editors. Methodologies for intelligent systems: 6th International Symposium, ISMIS ’91, Charlotte, N.C., USA, October 16–19, 1991: proceedings, volume 542 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1991. ISBN 3-540-54563-8 (Berlin), 0-387-54563-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 I582 1991.

Siegelmann:1996:NNS

[SNG96] Hava Siegelmann, Ephraim Nissan, and Alex Galperin. A novel neural/symbolic hybrid approach to optimized fuel-allocation and automated revision of heuristics in nuclear engineering. [with] appendix (E. nissan): A bird’s eye view of AI or ANN tools in nuclear engineering. Technical Report Report (c/o E. Nissan), CMS, Univ. of Greenwich, Woolwich, London, UK, 1996.

Tenenbaum:1992:PIC

[TRLA92] Gershon Tenenbaum, Tsily Raz-Liebermann, and Zvi Artzi, editors. Proceedings of the International Conference on Computer Applications in Sport and Physical Education, January 2– 6, 1992, Wingate Institute, Natania, Israel. Wingate Institute for Physical Education and Sport and the Zinman College of Physical Education, Natania, Israel, 1992. ISBN ???? LCCN GV436 .I576 1992.

Weiss:1996:GTHa

[wHW96a] Ephraim Nissan (with Hillel Weiss). For a generalized theory of hyperjoseph. part I. Technical Report Report (c/o E. Nissan), CMS, Univ. of Greenwich, Woolwich, London, UK, 1996.

Weiss:1996:GTHb

[wHW96b] Ephraim Nissan (with Hillel Weiss). For a generalized theory of hyperjoseph. part II. Technical Report Report (c/o E. Nissan), CMS, Univ. of Greenwich, Woolwich, London, UK, 1996.

Weiss:1996:GTHc

[wHW96c] Ephraim Nissan (with Hillel Weiss). For a generalized theory of hyperjoseph. part III. Technical Report Report (c/o E. Nissan), CMS, Univ. of Greenwich, Woolwich, London, UK,